OpenStudy (ttop0816):

Is my answer correct?? Evaluate x^2/y + x^2/z , if x = 10, y = -5, and z = -2. choices: 30 -30 -70 <-- i chose this! 70

OpenStudy (anonymous):

The equation is: \[\frac{x^2}{y}+\frac{x^2}{z}\] So we do: \[\left.\left.\left.\frac{x^2}{y}+\frac{x^2}{z}\right|_{x=10}\right|_{y=-5}\right|_{z=-2}\] \[=\left.\left.\frac{(10)^2}{y}+\frac{(10)^2}{z}\right|_{y=-5}\right|_{z=-2}\] \[=\left.\left.\frac{100}{y}+\frac{100}{z}\right|_{y=-5}\right|_{z=-2}\] \[=\left.\frac{100}{-5}+\frac{100}{z}\right|_{z=-2}\] \[=\left.-20+\frac{100}{z}\right|_{z=-2}\] \[=-20+\frac{100}{-2}\] \[=-20-50\] \[=-70\]

OpenStudy (anonymous):

You are correct :-)